How many moles of chlorine atoms are present in 3 moles of MgCl2  


1
Expert's answer
2021-02-03T07:23:32-0500

MgCl2"\to" 2Cl

"\nu" (Cl)= "\nu"(MgCl2)*2=3*2=6 mol